I have tried and tried to solve this! Can you help me?!
Mr. Graham invested some of his $18,000 in bonds that made a 5% profit and the rest in bonds that made a 12% profit. If the profit on the 12% bonds was $885 more than the profit on the 5% bonds, ow much did Mr. Graham invest in the 5% bonds?
I would really like to understand this!!

Answer by mananth(16949)

Mr. Graham invested some of his $18,000 in bonds that made a 5% profit and the rest in bonds that made a 12% profit. If the profit on the 12% bonds was $885 more than the profit on the 5% bonds, ow much did Mr. Graham invest in the 5% bonds?

Let him invest x in 5% bonds
5%bonds-----------------12%bonds-------------- Total
x------------------------18000-x---------------18000 ( Investment)

0.05x--------------------0.12(18000-x) ( interest)

The difference betwen 12% bond interest and 5% bond interest is $885
0.12(18000-x) -0.05x=885
2160 - 0.12x -0.05x = 885

-0.17x =-2160+885
-0.17x = -1275
x= -1275/0.17
=7500
He invested $ 7500 in 5% bonds